OpenStudy (anonymous):

show that: (1+cotx)/cscx - secx/(tanx + cotx) = cosx

OpenStudy (anonymous):

if you want a hint, i would rewrite everything without sines and cosines to make the algebra easier put \[\cos(x)=a,\sin(x)=b\]

OpenStudy (anonymous):

then you have \[\frac{1+\frac{a}{b}}{\frac{1}{b}}-\frac{\frac{1}{a}}{\frac{b}{a}+\frac{a}{b}}=a\]

OpenStudy (anonymous):

Input: 1 + cos(x) / csc(x) - sec(x) / tan(x) + cot(x) = cos(x) Results: True.

OpenStudy (anonymous):

do a bunch of algebra unencumbered by sines and cosines at some point you will probably need that \[\cos^2(x)+\sin^2(x)=1\] so if you see \[a^2+b^2\] replace it by 1 this is mostly just annoying algebra, there is little or no trig in it

OpenStudy (anonymous):

i would show the solution..the equation thingie is freaking weird..

OpenStudy (anonymous):

i can write it out for you if you like, that is i can write out the algebra

OpenStudy (anonymous):

...ok..............

OpenStudy (anonymous):

\[\frac{1+\frac{a}{b}}{\frac{1}{b}}-\frac{\frac{1}{a}}{\frac{b}{a}+\frac{a}{b}}=a\] multiply top and bottom of first fraction by b and second one by ab get \[b+a-\frac{b}{b^2+a^2}=a\] now replace \[a^2+b^2=1\] get \[b+a-b=a\] and you are done

OpenStudy (anonymous):

if you have to turn this in rewrite exactly what i wrote replacing a by cosine and b by sine

OpenStudy (callisto):

(1+cotx)/cscx - secx/(tanx + cotx) = 1/cscx + cotx/cscx - secx/ {[(sinx)^2 + (cosx)^2 ] / cosxsinx} = sinx + cotxsecx - secx / (1/sinxcosx) = sinx + (cosx/sinx)/(1/sinx) - secx(sinxcosx) =sinx + cosx - sinx = cosx